# A Deep Learning Method for Simultaneous Denoising and Missing Wedge Reconstruction in Cryogenic Electron Tomography

### v2

@@ 2.1 Denoising and missing wedge reconstruction algorithm (DeepDeWedge)
Step 3: Tomogram refinement. Pass the original, non-updated sub-tomograms $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\}_{i=1}^{N}$ through the fitted model $f_{\text{$\boldsymbol{\theta}$}^{*}}$ from Step 2, and reassemble the model outputs $\{f_{\text{$\boldsymbol{\theta}$}^{*}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\right)\}_{i=1}^{N}$ into a full-sized tomogram. Repeat the same for the sub-tomograms $\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{1}}}\}_{i=1}^{N}$. Finally, average both reconstructions to obtain the final denoised and missing-wedge corrected tomogram.
2.1.1 Motivation for the three steps of the algorithm
@@
Motivation for Step 1: We split the tilt series into two disjoint parts to obtain measurements with independent noise. As the noise on the individual projections or frames is assumed to be independent, the reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})$ are noisy observations of the same underlying sample with independent noise terms. Those are used in Step 2 for the self-supervised noise2noise-inspired loss.
@@
Tilt series splitting is also used in popular implementations of noise2noise-like denoising methods for cryo-ET [Buc+19, Bep+20, TC19]. The frame-based splitting procedure was proposed by [Buc+19, ], who found that it can improve the performance of noise2noise-like denoising over the even/odd split.
Motivation for Step 2: A brief justification for the loss function in Equation (2) is as follows; a detailed motivation is in Section 2.1.3.
@@
As the masks $\mathbf{M}\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}$, and $\mathbf{M}^{C}\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}$ are orthogonal, the loss value $\ell(i,\text{$\boldsymbol{\theta}$})$ can be expressed as the sum of two the two terms $\left|\left|\mathbf{M}\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}\mathbf{F}\left(f_{\text{$\boldsymbol{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,{\text{$\boldsymbol{\varphi}$}_{i}}}}^{\text{0}}}\right)-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}$, and $\left|\left|2\mathbf{M}^{C}\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}\mathbf{F}\left(f_{\text{$\boldsymbol{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}\right)-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}$. The first summand, i.e., $\left|\left|\mathbf{M}\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}\mathbf{F}\left(f_{\text{$\boldsymbol{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}\right)-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}$, is the squared L2 distance between the network output and the target sub-tomogram ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{1}}}$ on all Fourier components that were not masked out by the two missing wedge masks $\mathbf{M}$ and $\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}$. As we assume the noise in the target to be independent of the noise in the input ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}$, minimizing this part incentivizes the network to learn to denoise these Fourier components. This is inspired by the noise2noise principle.
@@
The second summand, i.e., $\left|\left|2\mathbf{M}^{C}\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}\mathbf{F}\left(f_{\text{$\boldsymbol{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}\right)-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}$, incentivizes the network $f_{\text{$\boldsymbol{\theta}$}}$ to restore the data that we artificially removed with the mask $\mathbf{M}$, and can be considered as a noisier2noise-like loss (see Section A.2 for background). For this part, it is important that we rotate both volumes, which moves their original missing wedges to a new, random location.
@@
In the last part of Step 2, we correct the missing wedges of the sub-tomograms $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\}_{i=1}^{N}$ using the current model $f_{\text{$\boldsymbol{\theta}$}}$. Therefore, as the model fitting proceeds, the model inputs $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}\}_{i=1}^{N}$ will more and more resemble sub-tomograms with only one missing wedge, i.e., the one we artificially remove from the partially corrected sub-tomograms $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\}_{i=1}^{N}$ with the mask $\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}.$ This is a heuristic which is supposed to help the model perform well on the original sub-tomograms $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\}_{i=1}^{N}$ and $\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{1}}}\}_{i=1}^{N}$, which have only one missing wedge and which we use as model inputs in Step 3. An analogous approach for noisier2noise-based image denoising was proposed by [Zha+22, ].
@@
Motivation for Step 3: To use all the information contained in the tilt series $\mathbf{t}$ for the final reconstruction, we separately refine the sub-tomograms from the FBP reconstructions $\text{FBP}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}}\right)$ and $\text{FBP}\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}}\right)$ of the sub-tilt-series ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}}$, and average them. Furthermore, in this step, we apply a special normalization to the tomograms, which is described in Appendix B.
2.1.2 Related methods

### v3

@@ 2 DeepDeWedge: Denoising and Missing Wedge Reconstruction Algorithm
    Tomogram refinement: Pass the original, non-updated sub-tomograms $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\}_{i=1}^{N}$ through the fitted model $f_{\text{$\bm{\theta}$}^{*}}$ from Step 2, and reassemble the model outputs $\{f_{\text{$\bm{\theta}$}^{*}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\right)\}_{i=1}^{N}$ into a full-sized tomogram. Repeat the same for the sub-tomograms $\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{1}}}\}_{i=1}^{N}$. Finally, average both reconstructions to obtain the final denoised and missing-wedge corrected tomogram.
2.1 Motivation for the Three Steps of the Algorithm
We now provide a brief motivation for each of the three steps of DeepDeWedge.
Motivation for Step 1.
@@ 2.1 Motivation for the Three Steps of the Algorithm
In Step 1, we split the tilt series into two disjoint parts to obtain measurements with independent noise. As the noise on the individual projections or frames is assumed to be independent, the reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})$ are noisy observations of the same underlying sample with independent noise terms. Those are used in Step 2 for the self-supervised Noise2Noise-inspired loss.
@@
Tilt series splitting is also used in popular implementations of Noise2Noise-like denoising methods for cryo-ET [11, 12, 13]. The frame-based splitting procedure was proposed by Buchholz et al. [11], who found that it can improve the performance of Noise2Noise-like denoising over the even/odd split.
Motivation for Step 2.
Step 2 of DeepDeWedge is to fit a neural network to perform denoising and missing wedge reconstruction, for which we have designed a specific loss function $\ell$ (Equation (2)). We provide a brief justification for the loss function here; a detailed theoretical motivation is presented in the following section.
@@
As the masks $\mathbf{M}\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}$, and $\mathbf{M}^{C}\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}$ are orthogonal, the loss value $\ell(i,\text{$\bm{\theta}$})$ can be expressed as the sum of two the two terms $\left|\left|\mathbf{M}\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}\mathbf{F}\left(\mathrm{f}_{\text{$\bm{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,{\text{$\bm{\varphi}$}_{i}}}}^{\text{0}}}\right)-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}$, and $\left|\left|2\mathbf{M}^{C}\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}\mathbf{F}\left(\mathrm{f}_{\text{$\bm{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}\right)-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}$. The first summand, i.e., $\left|\left|\mathbf{M}\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}\mathbf{F}\left(\mathrm{f}_{\text{$\bm{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}\right)-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}$, is the squared L2 distance between the network output and the target sub-tomogram ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{1}}}$ on all Fourier components that were not masked out by the two missing wedge masks $\mathbf{M}$ and $\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}$. As we assume the noise in the target to be independent of the noise in the input ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}$, minimizing this part incentivizes the network to learn to denoise these Fourier components. This is inspired by the Noise2Noise principle.
@@
The second summand, i.e., $\left|\left|2\mathbf{M}^{C}\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}\mathbf{F}\left(\mathrm{f}_{\text{$\bm{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}\right)-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}$, incentivizes the network $\mathrm{f}_{\text{$\bm{\theta}$}}$ to restore the data that we artificially removed with the mask $\mathbf{M}$, and can be considered as a Noisier2Noise-like loss [16] (see Supplementary Information 1.2 for background). For this part, it is important that we rotate both volumes, which moves their original missing wedges to a new, random location.
@@
In the last part of Step 2, we correct the missing wedges of the sub-tomograms $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\}_{i=1}^{N}$ using the current model $\mathrm{f}_{\text{$\bm{\theta}$}}$. Therefore, as the model fitting proceeds, the model inputs $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}\}_{i=1}^{N}$ will more and more resemble sub-tomograms with only one missing wedge, i.e., the one we artificially remove from the partially corrected sub-tomograms $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\}_{i=1}^{N}$ with the mask $\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}.$ This is a heuristic which is supposed to help the model perform well on the original sub-tomograms $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\}_{i=1}^{N}$ and $\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{1}}}\}_{i=1}^{N}$, which have only one missing wedge and which we use as model inputs in Step 3. An analogous approach for Noisier2Noise-based image denoising was proposed by Zhang et al. [17].
Motivation for Step 3.
@@
In Step 3, we use the fitted model from Step 2 to produce the final denoised and missing wedge filled tomogram. To use all the information contained in the tilt series $\mathbf{t}$ for the final reconstruction, we separately refine the sub-tomograms from the FBP reconstructions $\text{FBP}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}}\right)$ and $\text{FBP}\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}}\right)$ of the sub-tilt-series ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}}$, and average them. For this, we apply a special normalization to the model inputs, which is described in Supplementary Information 2.
2.2 Related Work